The angles in a triangle are such that the measure of one angle is twice the measure of the smallest angle, while the measure of the third angle is three times the measure of the smallest angle. Find the measures of all three angles, in degrees.
Found 2 solutions by ikleyn, mananth:
Answer by ikleyn(52781)

.

x + 2x + 3x = 180 degrees

    6x      = 180

     x      = 180/6 = 30  degrees.


ANSWER.  The angles are 30 degs, 60 degs and 90 degs.

Solved.



Answer by mananth(16946)

Let the smallest angle be x deg
the second angle will be 2x deg
The third will be 3x deg
x+2x+3x = 180 deg ( sum of angles of a triangle =180 deg)
6x =180 deg.
x = 30 deg
2x =60
3x =90
the other two angles 60 deg and 90 deg